Alkali Halides Nanotubes: Structure and Stability

Francisco Alberto Fernandez-Lima^{1}, Enio Frota da Silveira^{2}, Marco Antonio Chaer Nascimento

3

1Chemistry Department, Texas A&M University, College Station, TX, USA 2Physics Department, Pontifícia Universidade Católica, Rio de Janeiro, Brazil 3Chemistry Institute, Universidade Federal do Rio de Janeiro, Rio de Janeiro,

Brazil

In this talk we report the results of accurate quantum mechanical calculations on some alkali halide (LiF, NaCl, KBr) neutral clusters. The relative stability of (MX)_{n=2−28} structural isomers was studied using DFT (B3LYP/LAVCV3P**). For the members of the cubic and nanotube series Coupled Cluster calculations at the CCSD/SDD level were also performed. Møller-Plesset perturbation theory at the MP2/LAVCV3P** level of calculation was also used to check for possible inconsistencies in the DFT results for the (MX)_{n=1−10} clusters. Quantum dynamics calculations, at room temperature and atmospheric pressure, were also performed for the (LiF)_{28} octagon nanotube. A variety of structures has been found but in this talk we limit the discussion on the tube-like ones. Structural properties and possible mechanisms of formation will be discussed.

The alkali halide nanotubes present much smaller diameters, ranging from 0.241 nm to 0.67 nm (LiF, tetragonal-KBr hexagonal) when compared to the more traditional covalent nanotubes of C (1-2 nm), BN (1-3 nm) or GaN (30-200 nm). This is a consequence of the fact that the interactions are dominantly electrostatic and, therefore, much less strain is involved in making ring-type structures than in the covalent ones.

The results obtained for the (LiF)_{n} clusters show that overall the cubic series is highly stable as expected from structures which are debris of a crystalline bulk. Surprisingly, however, nanotube clusters with hexagonal and octagonal cross section present similar or higher stability than the corresponding cubic clusters. This result represents a new paradigm for 3D macro assembly of alkali halide clusters. This work was funded by CNPq, FAPERJ and Instituto Nacional de Materiais Complexos Funcionais. F.A.F-L acknowledges support from the National Institute of Health (Grant No. 1K99RR030188-01).